In 2008, it was reported that 10% of people at WCU are left handed. In 2020, wesystematically select 128 students and find that 11 are left handed. Test whether or not the number of people at WCU that are left handed has changed.

What is the test statistic? Round to 2 decimal places.
Answer #1

Solution:

Point estimate = sample proportion = = x / n = 0.086

Test statistics

z = ( - ) / *(1-) / n

= ( 0.086 - 0.10) / (0.10*0.90) / 128

= -0.53